A buffer solution of pH of 9.50 contains P mol of a weak base and 0.100 mol of a salt of its conjugate acid. When 100 cm3 of a 0.0600 mol.dm-3 solution of hydrochloric acid was added to this buffer solution the pH changes by 0.200 pH units. Use this information to calculate the value of Kb for the weak base.


1
Expert's answer
2021-10-28T07:48:15-0400

9.5+0.2= 9.7

9.7=(0.06+0.1)Kb

9.7= 0.1Kb

Kb = = 9.7/0.1

= 97